Project A has the following cash flows. Assume discount rate is 11.5%. Calculate the EAA for the project.

Year 0          1          2          3          4

Cash flow            -40       15        16        16        12

Group of answer choices

$15.29

$4.98

$5.63

$1.83
Answer #1

D.1.83

first let us know the present value of all the cash flows.

present value factor = 1 /(1+r)^n.

here,

r = 11.5%

=>0.115.

year cash flow PV factor cash flow * PV factor
0 -40 1/(1.115)^0=>1 -40*1=>-40
1 15 1/(1.115)^1=0.8969 (15*0.8969)=>13.4535
2 16

1/(1.115)^2=>0.8044

(16*0.8044)=>12.8704
3 16 1/(1.115)^3=>0.7214 (16*0.7214)=>11.5424
4 12 1/(1.115)^4=>0.6470 (12*0.6470)=>7.764
Net present value 5.63

EAA = npv / present value of annuity factor

here,

present value of annuity factor = [1- (1+r)^(-n)]/r

=>[1-(1.115)^(-4)]/0.115.

=>3.06961391.

EAA = $5.63/3.06961391

=>$1.83